Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#261 closed flight processing (fixed)

IPY09/11, flight day 215/2009, Helheim

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ipy/IPY09_11-2009_215_Helheim

Data arrived from ARSF via SATA disk Daphne03 on 24/08/09.

Scientific objective: Ice thickness measurements to add to a record of two major glaciers and to create a high res DEM.

Alpha 4-High

PI: T. Murray

Sensors:

  • Eagle (08/04/2010)
  • Hawk (08/04/2010)
  • Leica LIDAR (08/04/2010)
  • Digital Photographs (11/02/2010)

Change History (25)

comment:1 Changed 10 years ago by knpa

  • Priority changed from alpha 4 medium to alpha 4 high

comment:2 Changed 10 years ago by knpa

  • Description modified (diff)

comment:3 Changed 10 years ago by knpa

Corrected logsheet project code from GB09_21

comment:4 Changed 10 years ago by benj

  • Owner set to benj
  • Status changed from new to assigned

comment:5 Changed 10 years ago by benj

We've got Lidar data for three flightlines but IPAS for only the first two. Logsheet only has two lines recorded (and an IPAS problem mentioned) so presumably the third line was aborted and re-flown on one of the other days (218 or 219).

comment:6 Changed 10 years ago by benj

Turns out the first two lidar lines we have are test lines. Have downloaded the remaining three lines from Thelma (of which I think the last two are test lines again.

Also, the IPAS problem was because of a break in the data. We actually do have IPAS data for the whole flight.

comment:7 Changed 10 years ago by benj

The two lidar lines do appear to have an offset between them (order ~50cm in places, though it's hard to pin down). Unfortunately they only have a very small overlap region and that's at an angle (over a glacier, which obviously is neither flat nor smooth), which means I can't tell which of the potential error sources is causing it.

comment:8 Changed 10 years ago by mark1

Photography data tagged using log file since keep getting error that IPAS data does not overlap with photo time stamps.

comment:9 Changed 10 years ago by mark1

Photography delivery checked and ready to deliver.

comment:10 Changed 10 years ago by mark1

  • Description modified (diff)

Photographs dispatched to PI on 11/02/2010

comment:11 Changed 9 years ago by benj

  • Owner changed from benj to emca

comment:12 Changed 9 years ago by emca

LiDAR processing complete. Ready for delivery checking at
~airborne/workspace/IPY09_11-2009_215_Helheim/delivery/20100324/IPY09

comment:13 Changed 9 years ago by mark1

LIDAR delivery checked. All looks good. Ready to deliver.

comment:14 Changed 9 years ago by emca

Hyperspectral data ready for delivery checking at
~airborne/workspace/IPY09_11-2009_215_Helheim/delivery/20100407/IPY09_11

comment:15 Changed 9 years ago by mark1

Hyperspectral delivery checked, example commands tested and data looks good. Ready to deliver.

comment:16 Changed 9 years ago by emca

Line No. SCT
Eagle 10.04
Eagle 2-0.01
Hawk 10.00

Note: due to Hawk failure no data for Hawk line 2

comment:17 Changed 9 years ago by mark1

  • Description modified (diff)

Lidar and Hyperspectral data dispatched to PI on 08/04/2010

comment:18 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

LiDAR, Hyperspectral and camera rsync'd to
/users/rsg/arsf/arsf_data/2009/flight_data/ipy/IPY09_11-2009_215_Helheim

Ready for archiving

comment:19 Changed 9 years ago by knpa

  • Owner changed from emca to knpa

Started archiving.

comment:20 Changed 9 years ago by knpa

Lev3 files not present.

comment:21 Changed 9 years ago by knpa

no webcam data for this project

comment:22 Changed 9 years ago by knpa

Finished archiving prep, tarball ready to be downloaded by NEODC.

comment:23 Changed 9 years ago by knpa

Backed up by NEODC, transferred to archive disk.

comment:24 Changed 9 years ago by knpa

  • Milestone set to 2009 Data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

closing ticket

comment:25 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.